# Classification of nilpotent almost abelian Lie groups admitting left-invariant complex or symplectic structures

### v1

@@ 3. ${\rm(c)}$
If $2\leq n\leq m_{1}$, then $C^{n}=\begin{pmatrix}0&0\\ B^{n-1}v&B^{n}\end{pmatrix}$ with
$B^{n-1}v=\sum_{i=d}^{r}\,v_{i}B^{n-1}\epsilon_{1}^{i}+B^{n-1}w_{0}=\begin{cases}\begin{array}[]{cc}0,&\text{if }m_{d}\leq n-1\\ \sum_{i=d}^{r}\,v_{i}\epsilon_{n}^{i},&\text{if }m_{d}>n-1.\end{array}\end{cases}$
Hence, if $m_{d}\leq n-1$, then $B^{n-1}v=0$. On the other hand, if $m_{d}>n-1$, then $B^{n-1}v\neq 0$ and $B^{n-1}v\notin{\rm Im}\,B^{n}$. Therefore, we have to consider the following cases:
Case I.1: Assume that $m_{d}\leq n-1$ and $m_{1}\leq n-1$. In particular, $m_{i}<n$, for all $i=1,\dots,r$.
Under these conditions we have that $\dim(\operatorname{Ker}\,C^{n})=1+\dim(\operatorname{Ker}\,B^{n})$. By (2)
$\dim(\operatorname{Ker}\,B^{n})=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s=\sum_{i=1}^{r}m_{i}+s=m.$
So, $\dim(\operatorname{Ker}\,C^{n})=m+1$.
Now, since $m_{d}+1\leq n$, by (2)
$\dim(\operatorname{Ker}\,A^{n})=\sum_{i=1,i\neq d}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s=\sum_{i=1}^{r}{m_{i}}+1+s=m+1.$
Case I.2: Now, assume that $m_{d}\leq n-1$ and there exists $\ell<d$ such that $\ell=\max\{i:m_{i}>n-1\}$. In particular, $m_{i}<n$, for all $i=\ell+1,\dots,r$ and $m_{i}\geq n$, for all $i=1,\dots,\ell$.
Again, $\dim(\operatorname{Ker}\,C^{n})=1+\dim(\operatorname{Ker}\,B^{n})$. And, by (2)
$\dim(\operatorname{Ker}\,B^{n})=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s=\sum_{i=\ell+1}^{r}m_{i}+\ell n+s.$
So, $\dim(\operatorname{Ker}\,C^{n})=\sum_{i=\ell+1}^{r}m_{i}+\ell n+s+1$.
Now, since $m_{d}+1\leq n$, by (2)
$\dim(\operatorname{Ker}\,A^{n})=\sum_{i=1,i\neq d}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s=\sum_{i=\ell+1}^{r}m_{i}+\ell n+s+1.$
Case II.1: $m_{d}\ >n-1$ and there exists $\ell>d$ such that $\ell=\min\{i:m_{i}\leq n-1\}$. In particular, $m_{i}<n$, for all $i=\ell,\dots,r$ and $m_{i}\geq n$, for all $i=1,\dots,\ell-1$.
Since $B^{n-1}v\neq 0$ and $B^{n-1}v\notin{\rm Im}\,B^{n}$, then $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,B^{n})$. By (2)
$\dim(\operatorname{Ker}\,B^{n})=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s=\sum_{i=\ell}^{r}m_{i}+(\ell-1)n+s.$
Now, since $m_{d}+1>n$, by (2)
$\dim(\operatorname{Ker}\,A^{n})=\sum_{i=1,i\neq d}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s=\sum_{i=\ell}^{r}m_{i}+(\ell-1)n+s.$
Case II.2: $m_{d}>n-1$ and $m_{r}>n-1$. In particular, $m_{i}\geq n$, for all $i=1,\dots,r$.
Again, $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,B^{n})$ and by (2)
$\dim(\operatorname{Ker}\,B^{n})=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s=nr+s.$
Now, since $m_{d}+1>n$, by (2)
$\dim(\operatorname{Ker}\,A^{n})=\sum_{i=1,i\neq d}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s=nr+s.$
Therefore, $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,A^{n})$, for all $1\leq n\leq m_{1}$. Then, $C$ is conjugate to $A$.
Finally, we suppose that $v_{i}=0$, for all $i=1,\dots,r$. Since $v\notin{\rm Im}\,B$, we have $v=w_{0}\neq 0$.
Let $A=\mathcal{J}_{m_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}\oplus\mathcal{J}_{2}\oplus 0_{s-1}$. The matrices $C$ and $A$ are both nilpotent and $C^{m_{1}}=A^{m_{1}}=0$. We show next that $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,A^{n})$, for all $11\leq n\leq m_{1}-$ 1.
For $n=1$, we have $\dim(\operatorname{Ker}\,C)=\dim(\operatorname{Ker}\,B)=r+s=m$ and
$\dim(\operatorname{Ker}\,A)=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}})+\dim(\operatorname{Ker}\,\mathcal{J}_{2})+s-1=r+s=m.$
For $12\leq n\leq m_{1}-$ 1, since $B^{n-1}v=0$, we have
$\dim(\operatorname{Ker}\,C^{n})=1+\dim(\operatorname{Ker}\,B^{n})=1+\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s.$
@@
$\dim(\operatorname{Ker}\,A^{n})=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{2})+s-1=1+\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s.$
Therefore, $\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,A^{n})$, for all $11\leq n\leq m_{1}-$ 1. Then, $C$ is conjugate to $A$. ∎
@@
$B=\mathcal{J}_{m_{1}}^{\oplus q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{k}}^{\oplus q_{k}}\oplus 0_{s},\quad\;m_{1}>\cdots>m_{k}\geq 2,\;\;q_{i}\geq 1\text{ for all }i,\;\;s\geq 0.$
Then $C$ is conjugate to one and only one of the following matrices:
@@ 2. ${\rm(b)}$
    $\mathcal{J}_{m_{1}}^{\oplus q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{d-1}}^{\oplus q_{d-1}}\oplus\mathcal{J}_{m_{d}+1}\oplus\mathcal{J}_{m_{d}}^{\oplus q_{d}-1}\oplus\mathcal{J}_{m_{d+1}}^{\oplus q_{d+1}}\oplus\cdots\oplus\mathcal{J}_{m_{k}}^{\oplus q_{k}}\oplus 0_{s}$, for $d\in\{1,\dots,k\}$;
@@ 3. ${\rm(c)}$
    $\mathcal{J}_{m_{1}}^{\oplus q_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{k}}^{\oplus q_{k}}\oplus\mathcal{J}_{2}\oplus 0_{s-1}$, for $s>0$.
Given a positive integer $m$, we say that an $r$-tuple $\Lambda=(\lambda_{1},\ldots,\lambda_{r})$ is a partition of $m$ if $\lambda_{i}$ are positive integers such that $\lambda_{1}\geq\cdots\geq\lambda_{r}$ and $\sum_{i}\lambda_{i}=m$. It follows from the Jordan decomposition theorem that the conjugacy classes of nilpotent $m\times m$ matrices are in one-to-one correspondence with partitions of $m$. The correspondence is given as follows. Each non-zero nilpotent matrix $A$ is conjugate to a unique matrix of the form:
$\mathcal{J}_{n_{1}}^{\oplus p_{1}}\oplus\dots\oplus\mathcal{J}_{n_{k}}^{\oplus p_{k}}\oplus 0_{t},\quad k\geq 1,\;n_{1}>\dots>n_{k}\geq 2,\;p_{i}>0\;\text{forall }i,\;t\geq 0.$
We associate to $A$ the following partition $\Lambda(A)$ of $m$:
(3) $\Lambda(A):=(\underbrace{n_{1},\ldots,n_{1}}_{p_{1}},\ldots,\underbrace{n_{k},\ldots,n_{k}}_{p_{k}},\underbrace{1,\ldots,1}_{t}).$
The partition corresponding to the zero matrix $0_{m}$ is $(\underbrace{1,\ldots,1}_{m})$.
For a non-zero nilpotent matrix $A\in M(2n-1,{\mathbb{R}})$, we will provide necessary and sufficient conditions for the existence of complex or symplectic structures on $\mathfrak{g}_{A}$ in terms of $\Lambda(A)$ (see Theorems 3.10 and 4.10).
3. Classification of nilpotent almost abelian Lie algebras admitting complex structures
In this section, it will be useful to view $M(m,{\mathbb{C}})$ inside of $M(2m,{\mathbb{R}})$ as follows. Define $\varphi:M(m,{\mathbb{C}})\to M(2m,{\mathbb{R}})$ by:
@@ 3. Classification of nilpotent almost abelian Lie algebras admitting complex structures
Denoting by $I_{m}$ the $(m\times m)$-identity matrix and setting
@@
(6) $\begin{split}M_{J_{m}}(2m,{\mathbb{R}})&:=\left\{B\in M(2m,{\mathbb{R}}):BJ_{m}=J_{m}B\right\}\\ &=\left\{\begin{pmatrix}M_{1}&-M_{2}\\ M_{2}&M_{1}\end{pmatrix}:M_{1},M_{2}\in M(m,{\mathbb{R}})\right\}.\end{split}$
@@
We recall from [29, Lemma 6.1] the characterization of almost abelian Lie algebras with a complex structure (see also [2, Lemma 3.1]):

### v2

@@ 3. ${\rm(c)}$
For $2\leq n\leq m_{1}$, let $\ell=\max\{i\,:m_{i}\geq n\}$, i.e.,
$\text{either}\quad m_{1}\geq\dots\geq m_{\ell}\geq n>m_{\ell+1}\geq\cdots\geq m_{r},\quad\text{or }\;\;m_{r}\geq n.$
A straightforward computation shows that $C^{n}=\begin{pmatrix}0&0\\ B^{n-1}v&B^{n}\end{pmatrix}$ with
$B^{n-1}v=\sum_{i=d}^{r}\,v_{i}B^{n-1}\epsilon_{1}^{i}+B^{n-1}w_{0}=\begin{cases}\begin{array}[]{cc}0,&\text{if }n-1\geq m_{d}\\ \sum_{i=d}^{r}\,v_{i}\epsilon_{n}^{i},&\text{if }n-1<m_{d}.\end{array}\end{cases}$
Hence, we consider the following cases.
Case: $n-1\geq m_{d}$. Notice that in this case $\ell<d$, $\ell<r$ and $\dim(\operatorname{Ker}C^{n})=\dim(\operatorname{Ker}B^{n})+1$. We compute
$\dim(\ker B^{n})$ $=$ $\sum_{i=1}^{\ell}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+\sum_{i=\ell+1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+s=n\ell+\sum_{i=\ell+1}^{r}m_{i}+s.$
On the other hand, since $n\geq m_{d}+1$ and $d>\ell$, it follows that
$\dim(\ker A^{n})$ $=$ $\sum_{i=1}^{\ell}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+\sum_{\begin{smallmatrix}i=\ell+1\\ i\neq d\end{smallmatrix}}^{r}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s,$
$=$ $n\ell+\sum_{i=\ell+1}^{r}m_{i}+s+1=\dim(\operatorname{Ker}B^{n})+1=\dim(\operatorname{Ker}C^{n}).$
Case: $n-1<m_{d}$. Notice that, $d\leq\ell$ and $\dim(\ker C^{n})=\dim(\ker B^{n})$. Analogous to the previous case,
$\dim(\ker B^{n})=\begin{cases}n\ell+\sum_{i=\ell+1}^{r}m_{i}+s,&\text{if }\;\ell<r,\\ nr+s,&\text{if }\;\ell=r.\end{cases}$
Since $n<m_{d}+1$ and $d\leq\ell$, it follows that
$\dim(\ker A^{n})$ $=$ $\begin{cases}\sum_{\begin{smallmatrix}i=1\\ i\neq d\end{smallmatrix}}^{\ell}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+\sum_{i=\ell+1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s,&\text{if }\;\ell<r,\\ \sum_{\begin{smallmatrix}i=1\\ i\neq d\end{smallmatrix}}^{r}\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{i}})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{m_{d}+1})+s,&\text{if }\;\ell=r,\end{cases}$
$=$ $\begin{cases}n\ell+\sum_{i=\ell+1}^{r}m_{i}+s,&\text{if }\;\ell<r,\\ nr+s,&\text{if }\;\ell=r,\end{cases}$
$=$ $\dim(\operatorname{Ker}C^{n}).$
(c) If $v\notin{\rm Im}\,B\text{ and }v\in\text{span}\{\epsilon_{t+1},\ldots,\epsilon_{t+s}\}$. Let
$A=\mathcal{J}_{m_{1}}\oplus\cdots\oplus\mathcal{J}_{m_{r}}\oplus\mathcal{J}_{2}\oplus 0_{s-1}.$
In order to show that $C$ and $A$ are conjugate, it is enough to show that $\dim(\ker C^{n})=\dim(\ker A^{n})$, for all $11\leq n\leq m_{1}-$ 1, since $C^{m_{1}}=A^{m_{1}}=0$. For $n=1$, we have that $\dim(\operatorname{Ker}\,C)=\dim(\operatorname{Ker}\,B)=r+s$ and
$\dim(\operatorname{Ker}\,A)=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}})+\dim(\operatorname{Ker}\,\mathcal{J}_{2})+s-1=r+s.$
For $12\leq n\leq m_{1}-$ 1, since $B^{n-1}v=0$, it follows that
$\dim(\operatorname{Ker}\,C^{n})=\dim(\operatorname{Ker}\,B^{n})+1=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s+1.$
@@
$\dim(\operatorname{Ker}\,A^{n})=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+\dim(\operatorname{Ker}\,\mathcal{J}^{n}_{2})+s-1=\sum_{i=1}^{r}\dim(\operatorname{Ker}\,\mathcal{J}_{m_{i}}^{n})+s+1.$
Hence, $C$ and $A$ are conjugate.
Finally, if $B=0_{m}$ and $v\neq 0$, let $A=\mathcal{J}_{2}\oplus 0_{m-1}$, then $C^{2}=A^{2}=0$ and $\dim(\text{Ker}\,C)=m=\dim(\text{Ker}\,A)$. Therefore, $C$ is conjugate to $A$. ∎
@@
Let $C=\begin{pmatrix}0&0\\ v&B\end{pmatrix}\in M(m+1,{\mathbb{R}})$ be a nilpotent matrix, where $v\in\mathbb{R}^{m}$, $B\in M(m,{\mathbb{R}})$ with $\Lambda(B)=(m_{1},\ldots,m_{k};q_{1},\ldots,q_{k};s)$. Then $\Lambda(C)$ is one and only one of the following tuples:
@@ 2. ${\rm(b)}$
    $\begin{cases}(m_{1},\cdots,m_{d-1},m_{d},\cdots,m_{k};q_{1},\cdots,q_{d-1}+1,q_{d}-1,\cdots,q_{k};s),&\text{if }m_{d-1}=m_{d}+1,\\ \text{ or }&\\ (m_{1},\cdots,m_{d-1},m_{d}+1,m_{d},\cdots,m_{k};q_{1},\cdots,q_{d-1},1,q_{d}-1,\cdots,q_{k};s),&\text{if }m_{d-1}>m_{d}+1>m_{d};\end{cases}$ for $d\in\{1,\dots,k\}$;
@@ 3. ${\rm(c)}$
    $(m_{1},\ldots,m_{k},2;q_{1},\ldots,q_{k},1;s-1)$ for $s>0$.
If $B=0_{m}$ and $v\neq 0$, then $\Lambda(C)=(2;1;m-1)$.
3.2. The non-nilpotent case
The following results are analogous to those given in the previous section in the non-nilpotent setting.
Lemma 3.8.
Let $w\in{\mathbb{R}}^{k},$ $Q\in M(k,{\mathbb{R}})$ and $b\in{\mathbb{R}}$ such that $b\notin$ spec$(Q)$.
  1. (i)
    If $C=\left(\begin{array}[]{cc}b&0\\ w&Q\end{array}\right)$, then $C$ is conjugate to $\left(\begin{array}[]{cc}b&0\\ 0&Q\end{array}\right)$.
  2. (ii)
    If $C=\left(\begin{array}[]{c|c|c}b&0&0\\ \hline\cr u&R&0\\ \hline\cr w&0&Q\end{array}\right)$, with $u\in{\mathbb{R}}^{n}$, $R\in M_{b}(n,{\mathbb{R}})$, then $C$ is conjugate to $\left(\begin{array}[]{c|c|c}b&0&0\\ \hline\cr u&R&0\\ \hline\cr 0&0&Q\end{array}\right)$.
Proof.
Since $Q-bI$ is invertible, let $x=(Q-bI)^{-1}w$.
(i) We observe that $\left(\begin{array}[]{cc}b&0\\ 0&Q\end{array}\right)=P^{-1}CP$ with $P=\left(\begin{array}[]{cc}1&0\\ -x&I_{k}\end{array}\right)$.
(ii) By considering $P=\left(\begin{array}[]{c|c|c}1&0&0\\ \hline\cr 0&I_{n}&0\\ \hline\cr-x&0&I_{k}\end{array}\right)$, it is easy to see that $\left(\begin{array}[]{c|c|c}b&0&0\\ \hline\cr u&R&0\\ \hline\cr 0&0&Q\end{array}\right)=P^{-1}CP$. ∎
Given an almost abelian Lie algebra $\mathfrak{g}_{C}$ with $C\neq 0$, it follows from Proposition 2.1 that the family of Lie algebras isomorphic to $\mathfrak{g}_{C}$ is parametrized by the conjugacy classes of $C$ up to a nonzero multiple. The purpose of the next lemma is to choose an appropriate matrix $M$ which is conjugate to a multiple of $C$.
Lemma 3.9.
Let $b\in{\mathbb{R}}$, $R\in M_{b}(n,{\mathbb{R}})$, $u\in{\mathbb{R}}^{n}$ and $Q\in M(k,{\mathbb{R}})$ with $k\geq 1,\;b\notin$ spec$(Q)$. Consider
$C=\left(\begin{array}[]{c|c|c}b&0&0\\ \hline\cr u&R&0\\ \hline\cr 0&0&Q\end{array}\right).$
  1. (i)
    If $b\neq 0$, then $\mathfrak{g}_{C}$ is isomorphic to $\mathfrak{g}_{M}$ where
    $M=\left(\begin{array}[]{c|c|c}1&0&0\\ \hline\cr\dfrac{1}{b}\,u&\dfrac{1}{b}\,R&0\\ \hline\cr 0&0&\dfrac{1}{b}\,Q\end{array}\right).$
  2. (ii)
    If $b=0$, so that $R$ is nilpotent, and $\emph{spec}\,(Q)=\{\lambda_{1},\ldots,\lambda_{\ell}\}$ with $|\lambda_{1}|\geq\cdots\geq|\lambda_{\ell}|>0$, then $\mathfrak{g}_{C}$ is isomorphic to $\mathfrak{g}_{M}$ where $M$ is given by
    $M=\left(\begin{array}[]{c|c|c}0&0&0\\ \hline\cr\dfrac{1}{d}\,u&\dfrac{1}{d}\,R&0\\ \hline\cr 0&0&\dfrac{1}{d}\,Q\end{array}\right),\quad\text{where }d=|\lambda_{1}|.$
Moreover, if $\mathfrak{g}_{M}\cong\mathfrak{g}_{M_{1}}$ with
$M_{1}=\left(\begin{array}[]{c|c|c}0&0&0\\ \hline\cr u_{1}&R_{1}&0\\ \hline\cr 0&0&Q_{1}\end{array}\right),\;R_{1}\text{ nilpotent, }\emph{spec}\,(Q_{1})=\{\mu_{1},\ldots,\mu_{s}\},\;1=|\mu_{1}|\geq\cdots\geq|\mu_{s}|>0,$
then $s=\ell$ and $M_{1}$ is conjugate to ±$\pm M$.
Proof.
The proof of (i) and the first assertion in (ii) is a straightforward consequence of Proposition 2.1. To complete the proof of (ii), observe that
$\text{spec}\,\left(\frac{1}{d}\,Q\right)=\left\{\frac{\lambda_{1}}{d},\ldots,\frac{\lambda_{\ell}}{d}\right\},\quad\text{with}\quad 1=\frac{|\lambda_{1}|}{d}\geq\cdots\geq\frac{|\lambda_{\ell}|}{d}>0.$
Since $\mathfrak{g}_{M}$ is isomorphic to $\mathfrak{g}_{M_{1}}$, it follows from Proposition 2.1 that $M_{1}$ is conjugate to $cM$ for some non-zero real number $c$. In particular, $\text{spec}\,\left(Q_{1}\right)=\text{spec}\,\left(\dfrac{c}{d}\,Q\right)$. Therefore, $s=\ell$ and clearly
$\frac{|c|}{d}\,|\lambda_{1}|\geq\cdots\geq\frac{|c|}{d}\,|\lambda_{\ell}|,$
therefore, $1=|\mu_{1}|=\dfrac{|c|}{d}\,|\lambda_{1}|=|c|$, which implies $c=\pm 1$ since $c\in{\mathbb{R}}$. This completes the proof of (ii). ∎
4. Classification of almost abelian Lie algebras admitting complex structures
4.1. Characterization of almost abelian Lie algebras admitting a complex structure
In this section it will be useful to view $M(m,{\mathbb{C}})$ inside of $M(2m,{\mathbb{R}})$ as follows. Define $\varphi:M(m,{\mathbb{C}})\to M(2m,{\mathbb{R}})$ by:
@@ 4.1. Characterization of almost abelian Lie algebras admitting a complex structure
Setting
@@
$\begin{split}M_{J_{m}}(2m,{\mathbb{R}})&:=\left\{B\in M(2m,{\mathbb{R}}):BJ_{m}=J_{m}B\right\}\\ &=\left\{\begin{pmatrix}M_{1}&-M_{2}\\ M_{2}&M_{1}\end{pmatrix}:M_{1},M_{2}\in M(m,{\mathbb{R}})\right\}.\end{split}$
@@
We recall from [32, Lemma 6.1] the characterization of almost abelian Lie algebras with a complex structure (see also [3, Lemma 3.1]):
@@
Proposition 4.5.
Let $J\in M(2m,{\mathbb{R}})$ such that $J^{2}=-I_{2m}$ and let $B\in M(2m,{\mathbb{R}})$. Then $B$ is conjugate in $GL(2m,{\mathbb{R}})$ to a matrix in $M_{J}(2m,{\mathbb{R}})$ if and only if for any real eigenvalue $\lambda$ of $B$ the following condition is satisfied
(13) if $\text{if }\,\Lambda(B(\lambda))=\begin{cases}(n_{1},\ldots,n_{k};p_{1},\ldots,p_{k};t),&\text{ then }t\text{ and }p_{i}\text{ are even for all }i,\\ (m_{\lambda}),&\text{ then }\,m_{\lambda}\text{ is even}.\end{cases}$
In particular, if $B$ is conjugate to a matrix in $M_{J}(2m,{\mathbb{R}})$, then its semisimple and nilpotent parts are conjugate to matrices in $M_{J^{\prime}}(2m,{\mathbb{R}})$ and $M_{J^{\prime\prime}}(2m,{\mathbb{R}})$, respectively, for some $J^{\prime}$ and $J^{\prime\prime}$.
Proof.
Without loss of generality, we may suppose that $J=J_{1}^{\oplus m}$, where $J_{1}=\left(\begin{array}[]{cc}0&-1\\ 1&0\end{array}\right)$. Let $\psi$ be the injective ${\mathbb{R}}$-algebra homomorphism given by
$\psi:M(m,{\mathbb{C}})\rightarrow M(2m,{\mathbb{R}}),\hskip 14.22636pt\psi(A)=\left(\mathcal{C}(b_{kj},d_{kj})\right),\text{ for }A=(b_{kj}+id_{kj}),$
that is, $\psi(A)$ is the real $2m\times 2m$ matrix consisting of $m^{2}$ blocks of size $2\times 2$ of the form $\mathcal{C}(b_{kj},d_{kj})$ as in (3). It is straightforward that Im $\psi=M_{J}(2m,{\mathbb{R}})$. Also, it is easy to check that $\psi$ satisfies the following properties:
  1. (i)
    $\psi(A)$ is conjugate to $\psi(C)$ for conjugate matrices $A$, $C\in M(m,{\mathbb{C}})$;
  2. (ii)
    $\psi(A\oplus C)=\psi(A)\oplus\psi(C)$ for square matrices $A$ and $C$;
  3. (iii)
    for real numbers $a,b$ with $b>0$, $\psi(\mathcal{J}_{n}(a+ib))=\mathcal{C}_{n}(a,b)$ (see (4)) and $\psi(\mathcal{J}_{n}(a))$ is conjugate to $\mathcal{J}_{n}(a)\oplus\mathcal{J}_{n}(a)$.
Now, for $B\in M_{J}(2m,{\mathbb{R}})$, the matrix $\psi^{-1}(B)\in M(m,{\mathbb{C}})$ is conjugate to its Jordan normal form $\mathcal{J}(\psi^{-1}(B))$. Hence, by property (i) above, $B$ is conjugate to $\psi(\mathcal{J}(\psi^{-1}(B)))$. By conditions (ii) and (iii) we obtain the real Jordan normal form of $B$ (up to reordering). Finally, we have that the sets of distinct real eigenvalues of $\psi^{-1}(B)$ and $B$ coincide and if $\Lambda_{{\mathbb{C}}}(\psi^{-1}(B)(\lambda))=(m_{1},\ldots,m_{k};q_{1},\ldots,q_{k};s)$, where $\Lambda_{{\mathbb{C}}}$ is defined as in (2) for a matrix in $M_{\lambda}(r,{\mathbb{C}})$, property (iii) implies that $\Lambda(B(\lambda))=(m_{1},\ldots,m_{k};2q_{1},\ldots,2q_{k};2s)$ for any real eigenvalue $\lambda$. Therefore, condition (13) holds.
For the converse, suppose that $B\in M(2m,{\mathbb{R}})$ satisfies condition (13) and let
$\text{spec}(B)=\{\lambda_{1},\ldots,\lambda_{\ell},\lambda_{\ell+1},\overline{\lambda}_{\ell+1},\ldots,\lambda_{s},\overline{\lambda}_{s}\},$
where $\lambda_{j}\in{\mathbb{R}}$, for $j=1,\ldots,\ell$, and $\lambda_{j}\in{\mathbb{C}}$ with $\text{Im}(\lambda_{j})>0$, for $j=\ell+1,\ldots,s$. The matrix $B$ is conjugate to $B(\lambda_{1})\oplus\ldots\oplus B(\lambda_{s})$ (see (7)). By hypothesis, $B(\lambda_{j})\in M(m_{\lambda_{j}},{\mathbb{R}})$ with $m_{\lambda_{j}}=2r_{j}$, for any $j=1,\ldots,\ell$. Now, for each $j=\ell+1,\ldots,s$, let $n_{j}\in{\mathbb{N}}$ such that $B(\lambda_{j})\in M(2n_{j},{\mathbb{R}})$. By reordering the Jordan blocks in each $B(\lambda_{j})$, for all $j$ such that $\lambda_{j}\in{\mathbb{R}}$, we obtain that $B(\lambda_{1})\oplus\ldots\oplus B(\lambda_{s})$ is conjugate to a matrix in $M_{J}(2m,{\mathbb{R}})$ where $J=J_{r_{1}}\oplus\ldots\oplus J_{r_{\ell}}\oplus J_{1}^{\oplus n}$ (see (10) for the definition of the complex structures) with $n=n_{\ell+1}+\ldots+n_{s}$, and the converse follows.
Finally, decompose $B$ as $B=B_{s}+B_{n}$ with $B_{s}$ semisimple and $B_{n}$ nilpotent. It is easily verified that if $B$ satisfies condition (13), then $B_{n}$ and $B_{s}$ also satisfy this condition, and the last assertion follows. ∎
4.2. The nilpotent case